# LoXR: Performance Evaluation of Locally Executing LLMs on XR Devices

### v1

@@ 5.4 Results of Parallelization with BT and TT
Results of TT: Fig. 9 presents the TT results for thread counts of 1, 2, 4, 8, 16, and 32. Across all four devices, processing speed sees the most significant increase when increasing the thread count from 1 to 4. The speed gain continues at 8 threads (or shows minor degradation), but beyond 8 threads, performance begins to decline slightly, with further degradation beyond 16 threads. Apple Vision Pro (CPU) fails at thread counts of 16 and 32, indicating its limitations in achieving this level of parallelism.
@@ 5.5 Memory Consumption Results
Fig. 10 presents the memory consumption results for each model-device pair. The values represent the mean memory consumption from five experiments with varying batch sizes of 128, 256, 512, and 1024. As expected, memory consumption varies across models, with most showing consistent usage across different devices. Generally, within each model series, memory consumption increases with model size. However, minor exceptions exist, such as $m_{7}$ on Apple Vision Pro (GPU), which consumes more memory than other models in the Phi-3.1 series $[m_{6}-m_{11}]$. This size-based trend does not necessarily hold across series, even for models of similar sizes. For example, m2 (1.36 GB) consumes less memory than $m_{6}$ (1.32 GB) despite having a larger size, as they belong to different series. Similarly, $m_{7}$ (1.94 GB) requires more memory than $m_{5}$ (2.0 GB) and $m_{15}$ (2.05 GB), as all three belong to different series. Another example is $m_{12}$ (2.63 GB) consuming more memory than $m_{16}$ (2.81 GB), despite being smaller in size.
From a device-specific perspective, Apple Vision Pro demonstrates significantly lower memory consumption than the other devices, averaging $0.5782$ GB on the CPU and $0.5488$ GB on the GPU across all $17$ models. Among the remaining devices, Meta Quest 3 consumes the least memory at $2.3520$ GB, followed by Vivo X100 Pro at $2.3540$ GB, while Magic Leap 2 exhibits the highest at $2.3748$ GB. These findings suggest that Apple Vision Pro (both CPU and GPU) is approximately four times more memory-efficient than Magic Leap 2, Meta Quest 3, and Vivo X100 Pro. Its superior memory management and hardware optimizations make it the most efficient device in terms of memory consumption.
@@ 5.6 Battery Consumption Results
Figure 11 presents the battery consumption results over a 10-minute experiment. We observe that for the first two series, the Qwen Series and Vikhr-Gemma Series, battery consumption remains relatively low and with lower variation across all four devices. However, for larger models in the LLaMA-2 Series and Mistral-7B Series, battery consumption increases significantly. AVP shows a strong correlation between model size and battery consumption, both for GPU and CPU usage. In contrast, the remaining three devices do not exhibit a significant variation with respect to model size. Vivo demonstrates the best battery life, with an average loss of only $2.5\%$ over 10 minutes. ML2 follows with a $8.5\%$ average battery loss, while MQ3 records a slightly higher battery loss at $9.7\%$. AVP shows the worst in this regard, with GPU usage resulting in a $10.1\%$ battery loss and with CPU usage leading to a $12.6\%$ battery loss in 10 minuets.
Figure 11: Battery Consumption in 10-minute experiments.
@@ 5.7 Pareto Front Results
Since we use both CPU and GPU for inference in the Vision Pro experiments, while the other three devices use CPU only for inference, we exclude Apple Vision Pro GPU from the Pareto calculation to ensure a fair comparison. With 4 devices and 16 models (m2-m17), we have a total of 64 device-model pairs. Using Eq. 6 and Eq. 7, we calculated the quality, stability and performance scores for each pair, then identified the Pareto fronts and visualized them in Fig. 12. The quality of a model can be evaluated using various benchmarks, such as accuracy and perplexity, as detailed in Section 4.1. For the performance objective, we evaluate performance using PP, TG, memory consumption, and battery consumption, with respective weights of 0.35, 0.35, 0.2, and 0.1. For the stability objective, we assess it based on the coefficient of variation (CV) and error count of PP and TG, assigning 0.7 to CV and 0.3 to error count.
@@
Figure 12 illustrates the Pareto front points, highlighting the optimal choices across different model sizes and devices. For example, the AVP (CPU) device achieves the highest speed among the models but exhibits relatively low stability, whereas the ML 2 device offers the greatest stability. Meanwhile, both Vivo and Meta Q3 deliver a balanced trade-off between speed and stability. At the model-pair level, $m_{1}7$ and $m_{1}1$ provide high quality on AVP (CPU) and Vivo, respectively. At the same time, $m_{3}$ is selected as a Pareto-optimal choice due to its high performance on AVP (CPU) and robust stability on Meta Q3. A similar trend is observed with $m_{8}$, which attains strong performance and quality on AVP (CPU) as well as commendable quality and stability on ML 2.
@@ 6 Discussion
We have performed extensive testing of multiple open-source language models on four distinct devices, ranging from the two-year-old Magic Leap 2, representing the state of the art in passive see-through AR devices, to Meta Quest 3, arguably the most common VR headset worldwide, and VIVO X100s Pro as an indication of the performance of next-generation mobile computing, up to Apple Vision Pro, the highest-performance VR headset on the market. The performance in all devices shows that each of the devices currently available on the market is capable of LLM inference, achieving several tokens per second, and each of them can be used as an autonomous device for performing LLM inference for a given task.
The deployment of LLMs was achieved through a customized implementation of the Llama [22] library for each device. We selected Llama due to its C++-based architecture, which offers portability, flexibility, and seamless integration across various platforms and applications. This customization enables Llama to perform core LLMs tasks such as PP and TG, while providing the flexibility to adjust control parameters such as string length, batch size, and thread counts.
@@
Our study identifies critical performance evaluation factors and introduces a suite of testing scripts to analyze these aspects systematically. These scripts enhance the reproducibility of our findings, making them relevant for future analyses in similar contexts. The processing speed, measured in tokens per second, was evaluated across multiple dimensions. This includes a consistency analysis to determine whether each model-device pair maintains stable performance over time or exhibits fluctuations. The impact of string length on both PP and TG was analyzed for each model-device combination. Additionally, the study examined the support for parallelization and concurrency control through BT and TT tests. Furthermore, memory usage and battery consumption were also assessed to provide a holistic evaluation of system performance.
@@
For consistency analysis, we ran each model-device pair 20 times with conservative time intervals to observe performance variations. We reported the lowest and highest speeds, mean speed, standard deviation, and coefficient of variance (CV). Examining Figs. 4 and 3, which present the results in ascending order, we observe that the AVP GPU exhibits the highest stability. Among CPU-based devices, the AVP CPU and ML2 demonstrate the most stable performance. The quantitative results, particularly the CV values reported in Tables II and III, further reinforce this observation. While ML2 shows a similar level of stability to the AVP CPU, its first three models ($m_{2}$, $m_{3}$, and $m_{4}$) exhibit a CV above $20\%$ for TG and above $26\%$ for PP. Considering the error count in Fig. 7, ML2 emerges as the most stable device, having recorded zero errors. In contrast, the AVP CPU encountered multiple errors. The remaining two devices, Vivo and MQ3, exhibit higher instability, with Vivo performing relatively better than MQ3. This also indicates that consistency is model-dependent, as seen with models $m_{2}$, $m_{3}$, and $m_{4}$, which exhibited reduced stability.
@@
For $m_{1}$, being the smallest model, all devices performed efficiently. Among the remaining models ($m_{2}$ to $m_{16}$), the average inference speeds across various devices are as follows. AVP (CPU) achieved the highest inference speed (16.91 t/s PP, 11.04 t/s TG), followed by ML2 (12.68 t/s PP, 8.51 t/s TG), MQ3 (9.52 t/s PP, 5.77 t/s TG), and Vivo (8.62 t/s PP, 6.12 t/s TG), reflecting performance variations across XR devices. These speeds meet basic conversational requirements and enable LLM inference for broader applications beyond dialogue-based tasks. The PP speed across all devices is two to three times faster than TG speed, while for AVP GPU PP speed is 10 to 19 time faster than TG speed. This is because PP only encodes the input, whereas TG must process sequential dependencies during decoding. In model-based analysis, smaller models such as $m_{1}$, $m_{2}$, $m_{3}$, and $m_{5}$ are the fastest, while $m_{14}$ and $m_{16}$ were the slowest. Regarding string length, larger PP and TG inputs (512, 1024) are slower than shorter ones (64, 128, 256). Moreover, longer strings (512, 1024) frequently result in errors. Among devices, MQ3 exhibited the highest number of errors, followed by the AVP CPU and then Vivo. In contrast, ML2 and the AVP GPU did not experience any significant errors or inconsistencies.
@@
Battery consumption did not show significant differences across devices, although the Apple Vision Pro exhibited slightly higher battery life loss, with an average battery depletion of up to $10.1\%$ (with GPU) and $12.6\%$ (with CPU) over 10 minutes. In contrast, the Vivo X100 Pro performed the best, with only a $2.5\%$ loss, while the Magic Leap 2 demonstrated good battery efficiency with an $8.5\%$ loss. The Meta Quest 3 consumed an average of $9.7\%$ over the same duration.
@@
With 3D Pareto Optimality, we found AVP leading overall, ML2 demonstrating high consistency and stability, Vivo being slower but still consistent in performance, and MQ3 ranking lowest across all aspects. Among the models, $m_{3}$, and $m_{5}$ performed particularly well, with smaller models generally exhibiting better results. In contrast, models from the LLaMA-2 and Mistral-7B series were the slowest. Our study highlights the trade-offs between model complexity and system responsiveness. Larger models exhibit lower speed though they may offer greater accuracy and richer contextual responses. These larger models also drain batteries faster and decrease thermal stability. Conversely, lighter models enhance efficiency but may compromise response quality.
At this point, we can conclude that nowadays, the generation of XR devices is capable of executing interactive conversational tasks directly on the device. This is an important step towards XR device autonomy (and for edge devices in general) and an opportunity to perform complex tasks without any network connectivity requirement. We can anticipate that the next generation of XR devices with better computational capabilities, memory capacity, and battery life, will allow for complex on-device multimodal inference tasks, aiding users in a broad spectrum of tasks.

### v2

@@ 5.4 Results of Parallelization with BT and TT
Results of TT: Fig. 7 presents the TT results for thread counts of 1, 2, 4, 8, 16, and 32. Across all four devices, processing speed sees the most significant increase when increasing the thread count from 1 to 4. The speed gain continues at 8 threads (or shows minor degradation), but beyond 8 threads, performance begins to decline slightly, with further degradation beyond 16 threads. AVP (CPU) fails at thread counts of 16 and 32, indicating its limitations in achieving this level of parallelism.
@@ 5.5 Memory Consumption Results
Fig. 9 presents the memory consumption results for each model-device pair. The values represent the mean memory consumption from five experiments with varying batch sizes of 128, 256, 512, and 1024. As expected, memory consumption varies across models, with most showing consistent usage across different devices. Generally, within each model series, memory consumption increases with model size. However, minor exceptions exist, such as $m_{7}$ on AVP (GPU), which consumes more memory than other models in the Phi-3.1 series $[m_{6}-m_{11}]$. This size-based trend does not necessarily hold across series, even for models of similar sizes. For example, m2 (1.36 GB) consumes less memory than $m_{6}$ (1.32 GB) despite having a larger size, as they belong to different series. Similarly, $m_{7}$ (1.94 GB) requires more memory than $m_{5}$ (2.0 GB) and $m_{15}$ (2.05 GB), as all three belong to different series. Another example is $m_{12}$ (2.63 GB) consuming more memory than $m_{16}$ (2.81 GB), despite being smaller in size.
From a device-specific perspective, AVP demonstrates significantly lower memory consumption than the other devices, averaging $0.5782$ GB on the CPU and $0.5488$ GB on the GPU across all $17$ models. This efficiency is attributed to Apple Silicon’s unified memory architecture and the optimized memory management of the AVP, which can load only the active portions of a model during execution; for further details, see Feng et al. [57].
Among the remaining devices, Meta Quest 3 consumes the least memory at $2.3520$ GB, followed by Vivo X100 Pro at $2.3540$ GB, while Magic Leap 2 exhibits the highest at $2.3748$ GB. These findings suggest that AVP (both CPU and GPU) is approximately four times more memory-efficient than Magic Leap 2, Meta Quest 3, and Vivo X100 Pro. Its superior memory management and hardware optimizations make it the most efficient device in terms of memory consumption.
@@ 5.6 Battery Consumption Results
Fig. 10 presents the average battery consumption results over a 10-minute experiment, with each test repeated three times. We observe that for the first two series, the Qwen Series and Vikhr-Gemma Series, battery consumption remains relatively low and with lower variation across all four devices. However, for larger models in the LLaMA-2 Series and Mistral-7B Series, battery consumption increases significantly. AVP shows a strong correlation between model size and battery consumption, both for GPU and CPU usage. In contrast, the remaining three devices do not exhibit a significant variation with respect to model size. Vivo demonstrates the best battery life, with an average loss of only $2.5\%$ over 10 minutes. ML2 follows with a $8.5\%$ average battery loss, while MQ3 records a slightly higher battery loss at $9.7\%$. AVP shows the worst in this regard, with GPU usage resulting in a $10.1\%$ battery loss and with CPU usage leading to a $12.6\%$ battery loss in 10 minuets.
Figure 10: Battery consumption over 10-minute experiments, averaged across three runs.
@@ 5.7 Pareto Front Results
Since we use both CPU and GPU for inference in the Vision Pro experiments, while the other three devices use CPU only for inference, we exclude AVP GPU from the Pareto calculation to ensure a fair comparison. With 4 devices and 16 models (m2-m17), we have a total of 64 device-model pairs. Using Eq. 3 and Eq. 4, we calculated the quality, stability and performance scores for each pair, then identified the Pareto fronts and visualized them in Fig. 11. The quality of a model can be evaluated using various benchmarks, such as accuracy and perplexity, as detailed in Section 4.1. For the performance objective, we evaluate performance using PP, TG, memory consumption, and battery consumption, with respective weights of 0.35, 0.35, 0.2, and 0.1. For the stability objective, we assess it based on the coefficient of variation (CV) and error count of PP and TG, assigning 0.7 to CV and 0.3 to error count.
@@
Fig. 11 illustrates the Pareto front points, highlighting the optimal choices across different model sizes and devices. For example, the AVP (CPU) device achieves the highest speed among the models but exhibits relatively low stability, whereas the ML 2 device offers the greatest stability. Meanwhile, both Vivo and Meta Q3 deliver a balanced trade-off between speed and stability. At the model-pair level, $m_{1}7$ and $m_{1}1$ provide high quality on AVP (CPU) and Vivo, respectively. At the same time, $m_{3}$ is selected as a Pareto-optimal choice due to its high performance on AVP (CPU) and robust stability on Meta Q3. A similar trend is observed with $m_{8}$, which attains strong performance and quality on AVP (CPU) as well as commendable quality and stability on ML 2.
6 Analysis of On-Device LLMs
In this section, we evaluate the applicability of on-device LLMs in two interactive applications and report performance in terms of accuracy. Additionally, to assess efficiency, we compare on-device LLMs with alternative deployment approaches, namely client-server and cloud-based setups.
6.1 Evaluation on Interactive Applications
We conducted an analysis of our on-device LLM setup (AIvaluateXR) using datasets from two existing LLM-based interactive systems: VOICE [12] and GeoVis [58]. VOICE employs LLMs for interactive exploratory visualization and provides natural language interaction capabilities that allow users to navigate and manipulate 3D biological models in real time. GeoVis is another LLM-based method applied to geospatial data, enabling tasks such as selecting the viewing direction or retrieving location information through conversational queries. The details of the metrics, prompts and queries used are provided in the supplementary material. For this evaluation, we used the following metrics:
  • Formatting (GeoVis): Measures how well the model’s output conforms to the expected [latitude, longitude] format. A score of 100% is assigned if the formatting is perfect, 75% if it is off by one character, 50% if it is otherwise incorrect but still easy to parse, and 0% if it cannot be parsed.
  • Accuracy (GeoVis): Evaluates the spatial accuracy of the predicted coordinates relative to the target location. The score is 100% if the coordinates are near the center of the target area, 20% if they are off target but the area of interest remains partially visible at a reasonable zoom level, and 0% if it is not visible at all.
  • Score: For the VOICE dataset, each query is graded 100% if the output is correct and 0% otherwise. For the GeoVis dataset, if the Formatting score is 0, the overall Score is also 0. Otherwise, it is computed as a weighted combination of Formatting (20%) and Accuracy (80%) to reflect both structural and spatial correctness.
Since our current study focuses specifically on LLMs, we did not take evaluate the LLMs on prompts requiring a vision model or relying on fine-tuned models for modeling tasks. Instead, we used the VOICE and GeoVis datasets and evaluated only the queries that LLMs could answer on their own, which is sufficient for evaluating their accuracy. Table II presents the accuracy results. Since these tasks involve complex interactions, we observe that some models completely fail in terms of accuracy.
Table II: Performance summary of models $m_{8}$ to $m_{17}$ on the GeoVis [58] and VOICE [12] datasets. Each value $(\mu,\sigma)$ represents the mean ($\mu$) and standard deviation ($\sigma$), separated by a comma. The Score metric is computed as a weighted sum of Formatting (20%) and Accuracy (80%).
Model ID GeoVis data VOICE data
Formatting ($\mu,\sigma$) Accuracy ($\mu,\sigma$) Score ($\mu,\sigma$) Accuracy ($\mu,\sigma$)
$m_{8}$ 40.00 , 20.34 55.00 , 44.70 52.00 , 38.44 0.00 , 0.00
$m_{9}$ 40.00 , 20.34 57.33 , 46.23 51.20 , 39.85 0.00 , 0.00
$m_{10}$ 36.67 , 22.49 71.00 , 41.63 56.13 , 40.13 3.70 , 19.25
$m_{11}$ 45.00 , 15.26 64.00 , 45.00 60.20 , 37.57 0.00 , 0.00
$m_{12}$ 0.00 , 0.00 0.00 , 0.00 0.00 , 0.00 3.70 , 19.25
$m_{13}$ 1.67 , 9.13 0.00 , 0.00 0.33 , 1.83 0.00 , 0.00
$m_{14}$ 38.33 , 21.51 6.00 , 15.89 11.93 , 14.03 7.41 , 26.69
$m_{15}$ 52.50 , 7.63 39.33 , 42.74 41.97 , 34.92 48.15 , 50.92
$m_{16}$ 13.33 , 26.04 17.33 , 36.29 16.53 , 33.70 40.74 , 50.07
$m_{17}$ 54.17 , 13.27 87.33 , 29.47 80.70 , 23.63 51.85 , 50.92
GPT-4o 100 , 0 99.66 , 1.82 99.73 , 1.43 96.30 , 19.25
GPT-4o-Mini 98.33 , 9.13 93.33 , 22.94 94.33 , 19.79 85.19 , 36.20
Table III: Time comparison (in seconds), including both model loading (warm-up) time and first response latency, for model $M_{15}$ running locally on XR devices. Each cell reports the mean and standard deviation $(\mu,\sigma)$.
Model ID Series AVP ML2 Vivo MQ3
$m_{15}$ Mistral-7B 2.22 , 1.05 61.21 , 38.04 86.69 , 53.84 82.96,53.36
Table IV: Comparison of cold and warm first response time (F.R.T) in seconds for seven LLMs accessed via AVP across three deployment modes. Each cell reports the mean and standard deviation $(\mu,\sigma)$. Here, $\times$ indicates that the model is not supported in that mode.
Type Approach GPT-4o GPT-4o Mini Qwen2 ($m_{1}$) Vikhr-Gemma ($m_{5}$) Phi-3.1 ($m_{11}$) LLaMA-2 ($m_{13}$) Mistral ($m_{17}$)
Cold F.R.T Client-Server $\times$ $\times$ 0.67, 0.66 1.94, 1.36 2.22, 1.05 2.65, 1.54 3.72, 0.71
On-Device $\times$ $\times$ 1.07, 0.06 2.56, 0.37 3.06, 1.17 3.30, 0.93 3.86, 0.68
Warm F.R.T Cloud-based 1.13, 0.29 1.03, 0.31 $\times$ $\times$ $\times$ $\times$ $\times$
Client-Server $\times$ $\times$ 0.25, 0.15 1.36, 0.69 1.37, 0.07 1.40, 0.23 2.91, 0.68
On-Device $\times$ $\times$ 0.03, 0.00 0.26, 0.28 0.12, 0.04 0.27, 0.28 0.07, 0.00
6.2 Latency Comparison Across Deployment Modes
Latency, or first response time (F.R.T), refers to the duration between issuing a prompt and receiving the first token of the response, whereas warm-up time denotes the initial time required to load the model into memory. Cold F.R.T refers to the latency observed on the first prompt after model initialization (i.e., warm-up $+$ response), while Warm F.R.T reflects the prompt-to-first-token latency for subsequent queries on an already loaded model. We compared latency across three deployment modes: cloud-based, client-server, and on-device. In the client-server setup, the LLMs were hosted on a MacBook Pro with an M4 Pro chip and 48 GB of RAM, and accessed via AVP. To begin, we evaluated model $m_{15}$ across all four XR devices (see Table III) and selected the AVP as the fastest device for subsequent comparisons. For a broader evaluation, we selected five representative models (one per category) and benchmarked them across all three deployment modes. Each experiment was repeated three times, and we report the mean latency along with standard deviations.
As shown in Table IV, these five models were evaluated alongside GPT-4o and GPT-4o-mini, all accessed via the AVP. While on-device models exhibit longer warm-up times (i.e., higher Cold F.R.T), they consistently achieve the lowest latency after loading (i.e., lower Warm F.R.T), making them well-suited for real-time XR interactions, since warm-up only happens once during a session. Importantly, the latency values in Table IV correspond to a simple prompt—“Give me 20 continuous words as text only”. For longer or more complex prompts, the delay can further increase, potentially rendering cloud-based models unsuitable for time-sensitive XR applications.
[…]
Furthermore, over time, XR devices will feature more powerful hardware, and perhaps dedicated AI accelerators, so the latencies of on-device LLMs will diminish. On the other hand, cloud-based LLMs may gradually process prompts faster for the same reasons, but there will always be some network latency that is unlikely to shrink substantially. The relatively small difference in latency that we observed between GPT-4o and GPT-4o-mini suggests that network latency may already dominate overall latency for these models, indicating that this overall latency is unlikely to shrink by much [59]. Therefore, over time, the first-response time advantage that on-device LLMs currently enjoy is likely to grow, while their warm-up time disadvantage is likely to shrink, so long as the LLMs are small enough to be run efficiently on XR devices.
@@ 7 Discussion
AIvaluateXR provides a systematic framework for evaluating LLMs running locally on XR devices, taking into account both model- and device-specific factors. Using AIvaluateXR, we have performed extensive testing of multiple open-source language models on four distinct XR devices. Our study identifies critical performance evaluation factors and introduces a suite of testing scripts to analyze these aspects systematically. These scripts enhance the reproducibility of our findings, making them relevant for future analyses in similar contexts. The processing speed, measured in tokens per second, was evaluated across multiple dimensions. This includes a consistency analysis to determine whether each model-device pair maintains stable performance over time or exhibits fluctuations. The impact of string length on both PP and TG was analyzed for each model-device combination. Additionally, the study examined the support for parallelization and concurrency control through BT and TT tests. Furthermore, memory usage and battery consumption were also assessed to provide a holistic evaluation of system performance.
7.1 Summary of Results
@@
Figs. 4 and 3 present the consistency results in ascending order, where we see that the AVP GPU exhibits the highest stability. While ML2 shows a similar level of stability to the AVP CPU, its first three models ($m_{2}$, $m_{3}$, and $m_{4}$) exhibit a CV above $20\%$ for TG and above $26\%$ for PP. Considering the error count in Fig. 8, ML2 emerges as the most stable device, having recorded zero errors. In contrast, the AVP CPU encountered multiple errors. The remaining two devices, Vivo and MQ3, exhibit higher instability, with Vivo performing relatively better than MQ3. This also indicates that consistency is model-dependent, as seen with models $m_{2}$, $m_{3}$, and $m_{4}$, which exhibited reduced stability.
@@
For $m_{1}$, being the smallest model, all devices performed efficiently. Among the remaining models ($m_{2}$ to $m_{16}$), the average inference speeds across various devices (according to PP and TG tests) are as follows. AVP (CPU) achieved the highest inference speed (19.91 t/s PP, 11.54 t/s TG), followed by ML2 (12.18 t/s PP, 6.73 t/s TG), MQ3 (10.89 t/s PP, 5.17 t/s TG), and Vivo (9.35 t/s PP, 6.12 t/s TG), reflecting performance variations across XR devices. These speeds meet basic conversational requirements and enable LLM inference for broader applications beyond dialogue-based tasks. The PP speed across all devices is two to three times faster than TG speed, while for AVP GPU PP speed is 10 to 19 time faster than TG speed. This is because PP only encodes the input, whereas TG must process sequential dependencies during decoding. In model-based analysis, smaller models such as $m_{1}$, $m_{2}$, $m_{3}$, and $m_{5}$ are the fastest, while $m_{14}$ and $m_{16}$ were the slowest. Regarding string length, larger PP and TG inputs (512, 1024) are slower than shorter ones (64, 128, 256). Moreover, longer strings (512, 1024) frequently result in errors. Among devices, MQ3 exhibited the highest number of errors, followed by the AVP CPU and then Vivo. In contrast, ML2 and the AVP GPU did not experience any significant errors or inconsistencies.
@@
Battery consumption did not show significant differences across devices, although the AVP exhibited slightly higher battery life loss, with an average battery depletion of up to $10.1\%$ (with GPU) and $12.6\%$ (with CPU) over 10 minutes. In contrast, the Vivo X100 Pro performed the best, with only a $2.5\%$ loss, while the Magic Leap 2 demonstrated good battery efficiency with an $8.5\%$ loss. The Meta Quest 3 consumed an average of $9.7\%$ over the same duration.
@@
With 3D Pareto Optimality, we found AVP leading overall, ML2 demonstrating high consistency and stability, Vivo being slower but still consistent in performance, and MQ3 ranking lowest across all aspects. Among the models, $m_{3}$, and $m_{5}$ performed particularly well, with smaller models generally exhibiting better results. In contrast, models from the LLaMA-2 and Mistral-7B series were the slowest. We have analyzed the performance of an on-device LLM using 57 queries from an interactive application dataset. Our findings suggest that with targeted training and fine-tuning, the performance can be further improved.
7.2 The Impact of LLM Characteristics
The size of the model on disk, which correlates with the number of parameters, has a substantial impact on performance. In this study, we primarily selected models below 4 GB in size to ensure they could be deployed on all tested devices. We observed that larger models not only exhibited lower inference speeds and higher memory usage but also consumed more battery compared to smaller ones. However, in the context of interactive application accuracy, larger models such as the Mistral-7B and Phi-3.1 series achieved higher accuracy, while LLaMA-2 performed comparatively poorly. It is important to note that our testing for interactive applications was conducted without any fine-tuning. This suggests that selecting a smaller model—for example, from the Qwen series—and fine-tuning it on XR-specific data could achieve a better balance of accuracy and efficiency. Overall, these findings indicate the need to carefully balance model size and parameter count to reach optimal trade-offs between speed, memory consumption, battery life, and accuracy.
7.3 The Impact of Device Capabilities
We found that the AVP achieved the highest speed in both PP and TG tests. ML2 ranked second in overall performance, while the Vivo X100s Pro outperformed MQ3 in TG performance but ranked fourth in PP speed. These results can be attributed to the AVP’s unified memory architecture and hardware-accelerated memory management, which result in lower memory consumption and more stable performance. Furthermore, the combination of hardware design and operating system-level memory management plays a critical role in determining device efficiency. The remaining three devices demonstrated comparatively weaker performance, underscoring that selecting an XR device for AI workloads depends heavily on available memory and the underlying memory management capabilities. While understanding these factors beforehand is beneficial, applying AIvaluateXR provides a systematic way to quickly assess and compare device suitability for AI applications in XR environments.
7.4 Implications for XR Applications
Based on our evaluation of XR interactive applications, specifically VOICE and GeoVis (as shown in Table II), we observe that even without any domain-specific training, larger models achieved reasonably good accuracy. While smaller models demonstrated lower accuracy, their performance could likely be improved through grounding or fine-tuning on XR-specific datasets. The trade-offs between accuracy and processing time linked to model size are an important consideration for practical deployment but can be mitigated through targeted domain adaptation. These findings highlight the potential of AIvaluateXR as a practical tool for identifying and optimizing model-device configurations in this emerging area at the intersection of AI and XR.
7.5 The Future of On-Device LLMs
While cloud-based LLMs offer powerful capabilities, they may not always be preferred due to factors such as limited internet connectivity, subscription costs, privacy concerns, or regulatory restrictions. An alternative approach is the client-server model, where LLMs run on a local server and XR devices connect over a local network. However, this setup still introduces network latency (see Table IV) and does not fully address data security risks. By contrast, running LLMs directly on-device makes the system more autonomous and independent of external infrastructure. Looking ahead, we anticipate that advances in model compression, quantization techniques, and specialized hardware accelerators will further reduce resource demands and latency, making fully on-device intelligence increasingly practical. Given the rapid growth in XR and wearable devices, enabling on-device LLMs has the potential to unlock a wide range of real-world applications across domains, from education and training to healthcare, industrial maintenance, and beyond.
7.6 Limitations